# The Gorini-Kossakowski-Sudarshan-Lindblad problem and the geometry of CP maps

### v4

@@ 12 Conclusion
We have shown that the Jamiołkowski transform $\mathfrak{J}$ provides an excellent foundation on which to build the basic finite-dimensional theory of CP maps and of generators of Markovian CP evolutions in a unified way, and that the basic results can be extended to the separable setting by use of finite-dimensional approximations in the form of filtrations and their associated metrics. Among these basic results are: extreme CP maps are $\bm{\theta}$}-maps, every CP map has Kraus decompositions, a time-dependent element of $\mathsf{cp}^{+}(\mathcal{H})$ which is norm continuous generates a Markovian CP evolution, and Lindblad parametrizers exist. In addition, the classes $\mathsf{CP}(\mathcal{H},\mathcal{K})$, $\mathsf{cp}^{+}(\mathcal{H})$ and $\mathsf{cp}(\mathcal{H})$ are $d$-closed. It is significant that this extension can be done for Kraus decompositions, in particular, because there is an enormous gulf between the conventional approaches to the finite-dimensional and separable cases. Perhaps not so much as regards the GKSL problem, since Lindblad [2] does, in fact, make use of a sequence of finite-dimensional approximations in his key Proposition 5. However, there seems to be considerable difference in the way that primitive idea is used there compared to this paper.
@@ Appendix A Glossary/Index
$\mathcal{B}^{1}(\mathcal{H})$ trace-class operators, §7.1

### v5

@@ 12 Conclusion
We have shown that the Jamiołkowski transform $\mathfrak{J}$ provides an excellent foundation on which to build the basic finite-dimensional theory of CP maps and of generators of Markovian CP evolutions in a unified way, and that the basic results can be extended to the separable setting by use of finite-dimensional approximations in the form of filtrations and their associated metrics. Among these basic results are: extreme CP maps are $\bm{\theta}$}-maps, every CP map has Kraus decompositions, a time-dependent element of $\mathsf{cp}^{+}(\mathcal{H})$ which is norm continuous generates a Markovian CP evolution, and Lindblad parametrizers exist. In addition, the classes $\mathsf{CP}(\mathcal{H},\mathcal{K})$, $\mathsf{cp}^{+}(\mathcal{H})$ and $\mathsf{cp}(\mathcal{H})$ are $d$-closed. It is significant that this extension can be done for Kraus decompositions, in particular, because there is an enormous gulf between the conventional approaches to the finite-dimensional and separable cases. Perhaps not so much as regards the GKSL problem, since Lindblad [2] does, in fact, make use of a sequence of finite-dimensional approximations in his key Proposition 5. However, there seems to be considerable difference in the way that primitive idea is used there compared to this paper. As a practical matter, it is important to understand how finite-dimensional approximations work in the context of CP maps and Markovian evolutions. The approach of this paper makes a start on that project.
@@ Appendix A Glossary/Index
$\mathfrak{B}^{1}(\mathcal{H})$ trace-class operators, §7.1